For a populationof 17 year old boys and 17year old girls, the means and standard deviations, respectively, of their subscapular skinfold thickness values are as fllows : boys, 9.7 and 6.0: girls, 15.6 and 9.5. smple random samples of 40 boys and 35 girls are selected from the populations. What is the probability that the difference between sample means will be greater than 10?


1
Expert's answer
2021-06-21T11:00:15-0400

For boys-

Mean "\bar{x_1}=9.7,"

Standard deviation "\sigma_1=6"

"n_1=40"


For girls-

Mean "\bar{x_2}=15.6"

Standard deviation "\sigma_2=9.5"

"n_2=35"


Probability that the difference between sample means will be greater than 10-


"P(x_1-x_2>10)=P(Z>\dfrac{\bar{x_1}-\bar{x_2}}{\sqrt{\frac{\sigma_1^2}{n_1}+\frac{\sigma_2^2}{n_2}}})\\[9pt]=P(Z>\dfrac{9.7-15.6}{\sqrt{\frac{6^2}{40}+\frac{9.5^2}{35}}})=P(Z>-3.146)=0.9917"